ravi is 40 years older than tony. in 5 years, ravi would be twice the age of tony. the current ageof ravi and tony

Answer by stanbon(75887)

ravi is 40 years older than tony. in 5 years, ravi would be twice the age of tony. the current age of ravi and tony.
-----
Equations:
r = t + 40
r + 5 = 2(t + 5)
-----------
Substitute for "f" and solve for "t"::
t+40 + 5 = 2t + 10
t = 35 (tony's age now)
----
Solve for "r"::
r = t + 40 = 75 (ravi's age now)
-----------
Cheers,
Stan H.
-----------